OpenStudy (anonymous):

Frank and Gertrude mow lawns for extra money over the summer. Frank's income is determined by f(x) = 5x + 15, where x is the number of hours. Gertrude's income is g(x) = 4x + 20. If Frank and Gertrude were to combine their efforts, their income would be h(x) = f(x) + g(x). Assume Frank works 4 hours. Create the function h(x) and indicate if Frank will make more money working alone or by teaming with Gertrude.

OpenStudy (anonymous):

h(x) = x + 5, work alone h(x) = x + 5, team with Gertrude h(x) = 9x + 35, work alone h(x) = 9x + 35, team with Gertrude